# Robust Weight Initialization for Tanh Neural Networks with Fixed Point Analysis

### v1

@@ 3.3 Preventing activation saturation via appropriate $\sigma_{z}$ tuning
In this subsection, we discuss how $\sigma_{z}$ impacts the scale of the activation values. Equation 2 indicates that $a_{i}^{k}$ follows a normal distribution, with variance depending on $\sigma_{z}$. Firstly, we experimentally investigated the impact of $\sigma_{z}$ on the scale of the activation values. As demonstrated in Figure 2, increasing $\sigma_{z}=\alpha/\sqrt{N_{\ell-1}}$ causes the activation values in any layer to be distributed over a broader range. However, setting $\sigma_{z}$ to a large value can lead to saturation, where most activations converge towards $-1$ and $1$. If $\sigma_{z}$ is too large, the probability that $a_{i}^{(k)}$ takes values far from $1$ (e.g., -10, 5, etc.) increases. This, in turn, increases the value of $1+\epsilon$ mentioned in Corollary 4, potentially bounding the activation values in sufficiently deep layers by $\xi_{1+\epsilon}$. Consequently, the activation values in deeper layers become less likely to approach zero and tend to saturate toward specific values. On the other hand, if $\sigma_{z}$ is too small, as mentioned in Subsection 3.2, the variance of $a_{i}^{k}$ becomes restricted. This is demonstrated experimentally in Figure 2, when $\alpha=0.00001$. For this reason, we experimentally found an optimal $\sigma_{z}=\alpha/\sqrt{N^{\ell-1}}$, with $\alpha=0.085$, that is neither too large nor too small. Results from experiments solving the Burgers’ equation using PINNs with varying $\sigma_{z}$ are presented in Appendix B.1.
@@
Table 1: Validation accuracy and loss are presented for FFNNs with varying numbers of nodes $(2,8,32,128)$, each with $20$ hidden layers using the tanh activation function. All models were trained for $20$ epochs, and the highest average accuracy and lowest average loss, computed from $10$ runs, are presented. When comparing different initialization methods under the same experimental settings, the better-performing method is highlighted in bold. Underlined values indicate the highest accuracy when only the number of nodes is varied.
2 8 32 128
MNIST Accuracy Loss Accuracy Loss Accuracy Loss Accuracy Loss
Xavier 49.78 1.632 68 0.958 91.67 0.277 95.45 0.154
Proposed 62.82 1.185 77.95 0.706 92.51 0.255 96.12 0.134
FMNIST Accuracy Loss Accuracy Loss Accuracy Loss Accuracy Loss
Xavier 42.89 1.559 68.55 0.890 81.03 0.533 86.2 0.389
Proposed 51.65 1.324 71.31 0.777 83.06 0.475 87.12 0.359
CIFAR10 Accuracy Loss Accuracy Loss Accuracy Loss Accuracy Loss
Xavier 32.82 1.921 43.51 1.608 48.62 1.473 47.58 1.510
Proposed 38.16 1.780 47.04 1.505 48.80 1.463 48.51 1.471
@@
Table 2: Validation accuracy and loss are presented for FFNNs with varying numbers of layers $(10,50,100)$, each with $64$ number of nodes using the tanh activation function. All models were trained for $40$ epochs, and the highest average accuracy and lowest average loss, computed from $10$ runs, are presented.
10 50 100
MNIST Accuracy Loss Accuracy Loss Accuracy Loss
Xavier 96.55 0.112 96.57 0.123 94.08 0.194
Proposed 97.04 0.102 96.72 0.109 96.06 0.132
FMNIST Accuracy Loss Accuracy Loss Accuracy Loss
Xavier 88.73 0.319 87.72 0.344 83.41 0.463
Proposed 89.42 0.305 88.51 0.324 86.01 0.382
CIFAR10 Accuracy Loss Accuracy Loss Accuracy Loss
Xavier 48.39 1.468 47.87 1.474 46.71 1.503
Proposed 48.41 1.458 48.71 1.461 48.96 1.437
@@ 4 Experiments
In this section, we conduct a series of experiments to validate the proposed weight initialization method. In Subsection 4.1, we evaluate the performance of an FFNN with the tanh activation function on benchmark datasets. In Subsection 4.2, we solve the Burgers’ equation and Allen-Cahn equation using Physics-Informed Neural Networks. Both experiments are conducted across various network sizes to verify whether the proposed method consistently performs well, independent of network depth and width. The experiments were conducted in TensorFlow without skip connections, normalization layers, and learning rate decay in any of the experiments.
4.1 Width Independence in Classification Task
@@
To evaluate the effectiveness of the proposed weight initialization method, we conduct experiments on the MNIST, Fashion MNIST, and CIFAR-10 (Krizhevsky & Hinton, 2009) datasets, utilizing the Adam optimizer. All experiments are conducted with a batch size of 64 and a learning rate of 0.0001. Fifteen percent of the total dataset is allocated for validation.
@@
We apply the proposed weight initialization method to evaluate its effectiveness in training tanh FFNNs, emphasizing its robustness to variations in network width. Four tanh FFNNs are created, each with 20 hidden layers, and with 2, 8, 32, and 128 nodes per hidden layer, respectively. In Table 1, for both the MNIST and Fashion MNIST datasets, the network with 128 nodes achieves the highest accuracy and lowest loss when our proposed method is employed. However, for the CIFAR-10 dataset, the network with 32 nodes yields the highest accuracy and lowest loss when employing the proposed method. In summary, our proposed method demonstrates robustness regardless of the number of nodes in tanh FFNNs. We provide more detailed experimental results in Appendix B.2.
4.2 Depth Independence in Classification Task
@@
It is well known that the expressivity of neural networks generally increases exponentially with depth, enabling strong generalization performance (Poole et al., 2016; Raghu et al., 2017). Therefore, we employ the proposed weight initialization method to investigate its effectiveness in training deep FFNNs with the tanh activation function, emphasizing its robustness to variations in network depth. We create three tanh FFNNs, each with 64 nodes in all hidden layers, but with 10, 50, and 100 hidden layers, respectively. In Table 2, for both the MNIST and Fashion MNIST datasets, the network with 10 hidden layers achieves the highest accuracy and lowest loss when our proposed method is employed. Both initialization methods perform best in networks with the fewest layers, with performance degrading as the depth increases. However, for the CIFAR-10 dataset, we observe that the performance of the proposed method improves as the number of layers increases.
@@
Furthermore, we conduct experiments with varying hidden layer dimensions, as shown in Figure 3. The network consists of 60 hidden layers, where the number of nodes alternates between 32 and 16 in each layer. We demonstrate superior performance in terms of both loss and accuracy across all epochs on the MNIST and CIFAR-10 datasets.
@@
4.3 Network size independence in pinn
Xavier initialization is the primary method used for training PINNs  (Jin et al., 2021; Son et al., 2023; Yao et al., 2023; Gnanasambandam et al., 2023). In this section, we experimentally demonstrate that the method’s training performance is highly dependent on randomness and network size. Additionally, empirical results are provided demonstrating that the proposed method is more robust to variations in network size.
@@ 4.3 Network size independence in pinn
All experiments on Physics-Informed Neural Networks (PINNs) use full-batch training with a learning rate of 0.001. In this section, we solve the Allen-Cahn and Burgers’ equations using a tanh FFNN-based PINN with 20,000 collocation points. For the Allen-Cahn equation, the diffusion coefficient is set to $d=0.01$. The initial condition is defined as $u(x,0)=x^{2}\cos(\pi x)$ for $x\in[-1,1]$, with boundary conditions $u(-1,t)=-1$ and $u(1,t)=-1$, applied over the time interval $t\in[0,1]$. Similarly, for the Burgers’ equation, a viscosity coefficient of $\nu=0.01$ is employed. The initial condition is given by $u(x,0)=-\sin(\pi x)$ for $x\in[-1,1]$, with boundary conditions $u(-1,t)=0$ and $u(1,t)=0$ imposed for $t\in[0,1]$.
The Allen-Cahn equation is expressed as:
$\frac{\partial u}{\partial t}-d\frac{\partial^{2}u}{\partial x^{2}}=-\frac{u^{3}+u}{d}$
The Burgers’ equation is given by:
$\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}=\nu\frac{\partial^{2}u}{\partial x^{2}}$
Eight tanh FFNNs are created, each with 16 nodes in all hidden layers, but with 5, 10, 20, 30, 40, 50, 60, and 80 hidden layers, respectively. As shown in Table 3, for the Allen-Cahn equation, Xavier initialization achieves the lowest loss at a network depth of 20. However, as the depth increases, the loss gradually rises. In contrast, the proposed method achieves the lowest loss at a depth of 50 and maintains a loss of 0.00057 even at a depth of 80 layers. For the Burgers’ equation, the proposed method achieves the lowest loss at a depth of 60, while at the same depth, Xavier initialization results in a loss that is an order of magnitude higher (approximately $10^{2}$ difference).
Next, we double the number of nodes to observe the impact of node size on the loss. Eight new tanh FFNNs are created, each with 32 nodes in all hidden layers, and with 5, 10, 20, 30, 40, 50, 60, and 80 hidden layers, respectively. As shown in Table 3, for the Allen-Cahn equation, Xavier initialization achieves the lowest loss at a network depth of 30. However, as the depth increases, the model becomes untrainable, with a loss of 0.694 at a depth of 80. In contrast, the proposed method achieves the lowest loss at a depth of 40 and maintains a loss of 0.00059 even at a depth of 80 layers. For the Burgers’ equation, both methods show similar loss values up to a depth of 30. Beyond a depth of 40, however, the loss steadily increases with the Xavier method, while the proposed method records the lowest loss at a depth of 50.
@@
Table 3: A PINN loss is presented for FFNNs with varying numbers of layers $(5,10,20,30,40,50,60,80)$ using the tanh activation function. The top table shows results with $16$ nodes per layer, and the bottom table shows results with $32$ nodes per layer. All models were trained for $300$ iterations using Adam and $300$ iterations using L-BFGS. The median PINN loss from the final iteration for the Burgers and Allen–Cahn equations, computed over 5 runs, is presented.

### v2

@@ 3.3 Preventing activation saturation via appropriate $\sigma_{z}$ tuning
In this section, we determine the appropriate value of $\alpha$ in $\sigma_{z}=\alpha/\sqrt{N_{\ell-1}}$ that satisfies condition ((((ii)))). Condition ((((ii)))) is motivated by normalization methods (Ioffe, 2015; Ba, 2016). Firstly, we experimentally investigated the impact of $\sigma_{z}$ on the scale of the activation values. As shown in Figure 2, increasing $\sigma_{z}=\alpha/\sqrt{N_{\ell-1}}$ broadens the activation range in each layer, while decreasing $\sigma_{z}$ narrows it.
When $\sigma_{z}$ is Large. Setting $\sigma_{z}$ to a large value can lead to saturation. If $\sigma_{z}$ is too large, Equation (2) implies that the likelihood of $a_{i}^{k}$ deviating significantly from 1 increases. This increases the likelihood of activation values being bounded by $\xi_{1+\epsilon}$ in sufficiently deep layers, as stated in Corollary 4. Consequently, in deeper layers, activation values are less likely to approach zero and tend to saturate toward specific values. Please refer to the Figure 3 for the cases where $\sigma_{z}=0.3$ and $3$.
Figure 3: The activation values in the 1000th layer, with $32$ nodes per hidden layer, were analyzed using the proposed weight initialization method with $\sigma_{z}$ values of $0.0003$, $0.015$, $0.3$, and $3$. The analysis was conducted on $3,000$ input samples uniformly distributed within the range $[-1,1]$.
When $\sigma_{z}$ is Small. If $\sigma_{z}$ is too small, Equation (2) implies that the distribution of $a_{i}^{k}$ has a standard deviation close to zero. Consequently, $x_{i}^{k+1}$ can be approximated as the result of applying $\tanh(x)$ to $x_{i}>0$ repeatedly for a finite number of iterations, $k$. Since $\tanh^{\prime}(x)$ decreases for $x\geq 0$, the values resulting from finite iterations eventually saturate. Plese refer to the Figure 3 when $\sigma_{z}=0.0003$.
For these reasons, we experimentally determined an optimal $\sigma_{z}$ that avoids being excessively large or small. As shown in Figure 3, $\sigma_{z}=0.015$ maintains an approximately normal activation distribution without collapse. Additional experimental results are provided in Appendix A.5. Considering the number of hidden layer nodes, we set $\sigma_{z}=\alpha/\sqrt{N^{\ell-1}}$ with $\alpha=0.085$. Experimental results for solving the Burgers’ equation using PINNs with varying $\sigma_{z}$ are provided in Appendix C.3.
@@
Table 1: Validation accuracy and loss are presented for FFNNs with varying numbers of nodes $(2,8,32,128,512)$, each with $20$ hidden layers using tanh activation function. All models were trained for $20$ epochs, and the highest average accuracy and lowest average loss, computed over $10$ runs, are presented. The better-performing method is highlighted in bold when comparing different initialization methods under the same experimental settings.
Dataset Method 2 Nodes 8 Nodes 32 Nodes 128 Nodes 512 Nodes
Accuracy Loss Accuracy Loss Accuracy Loss Accuracy Loss Accuracy Loss
MNIST Xavier 49.78 1.632 68 0.958 91.67 0.277 95.45 0.154 97.35 0.087
Proposed 62.82 1.185 77.95 0.706 92.51 0.255 96.12 0.134 97.96 0.067
FMNIST Xavier 42.89 1.559 68.55 0.890 81.03 0.533 86.20 0.389 88.28 0.331
Proposed 51.65 1.324 71.31 0.777 83.06 0.475 87.12 0.359 88.59 0.323
CIFAR-10 Xavier 32.82 1.921 43.51 1.608 48.62 1.473 47.58 1.510 51.71 1.369
Proposed 38.16 1.780 47.04 1.505 48.80 1.463 48.51 1.471 52.21 1.359
CIFAR-100 Xavier 10.87 4.065 18.53 3.619 23.71 3.301 23.83 3.324 17.72 3.672
Proposed 15.22 3.818 23.07 3.350 24.93 3.237 24.91 3.240 22.80 3.435
@@
Table 2: Validation accuracy and loss are presented for FFNNs with varying numbers of layers $(3,10,50,100)$, each with $64$ number of nodes using the tanh activation function. All models were trained for $40$ epochs, and the highest average accuracy and lowest average loss, computed over $10$ runs, are presented.
Dataset Method 3 Layers 10 Layers 50 Layers 100 Layers
Accuracy Loss Accuracy Loss Accuracy Loss Accuracy Loss
MNIST Xavier 95.98 0.130 96.55 0.112 96.57 0.123 94.08 0.194
Proposed 96.32 0.123 97.04 0.102 96.72 0.109 96.06 0.132
FMNIST Xavier 85.91 0.401 88.73 0.319 87.72 0.344 83.41 0.463
Proposed 86.51 0.379 89.42 0.305 88.51 0.324 86.01 0.382
CIFAR-10 Xavier 42.91 1.643 48.39 1.468 47.87 1.474 46.71 1.503
Proposed 45.05 1.588 48.41 1.458 48.71 1.461 48.96 1.437
CIFAR-100 Xavier 19.10 3.628 22.73 3.400 24.27 3.283 20.32 3.515
Proposed 19.30 3.609 23.83 3.309 25.07 3.190 24.41 3.234
@@ 4 Experiments
In this section, we present a series of experiments to evaluate the proposed weight initialization method. In Section 4.1, we evaluate the performance of tanh FFNNs on benchmark classification datasets. In Section 4.2, we apply Physics-Informed Neural Networks to solve PDEs. Both experiments assess the proposed method’s robustness to network size and data efficiency.
4.1 Classification Task
@@
Experimental Setting. To evaluate the effectiveness of the proposed weight initialization method, we conduct experiments on the MNIST, Fashion MNIST (FMNIST), CIFAR-10, and CIFAR-100 (Krizhevsky & Hinton, 2009) datasets with the Adam optimizer. All experiments are performed with a batch size of $64$ and a learning rate of $0.0001$. Fifteen percent of the dataset is allocated for validation. The experiments are implemented in TensorFlow without skip connections and learning rate decay in any of the experiments.
@@
Width Independence in Classification Task. We evaluate the proposed weight initialization method in training tanh FFNNs, focusing on its robustness across different network widths. Five tanh FFNNs are designed, each with 20 hidden layers, and with $2,8,32$, $128$, and $512$ nodes per hidden layer, respectively. In Table 1, for the MNIST, Fashion MNIST and CIFAR-10 datasets, the network with 512 nodes achieves the highest accuracy and lowest loss when the proposed method is employed. However, for the CIFAR-100 dataset, the network with 32 nodes yields the highest accuracy and lowest loss when employing the proposed method. In summary, the proposed method demonstrates robustness across different network widths in tanh FFNNs. Detailed experimental results are provided in Appendix B.1.
@@
Depth Independence in Classification Task. The expressivity of neural networks is known to increase exponentially with depth, leading to improved generalization performance (Poole et al., 2016; Raghu et al., 2017). To evaluate the robustness of the proposed weight initialization method across different network depths, we conduct experiments on deep FFNNs with tanh activation functions. Specifically, we construct four tanh FFNNs, each with $64$ nodes per hidden layer and $3$, $10$, $50$, and $100$ hidden layers. respectively. In Table 2, for both the MNIST and Fashion MNIST datasets, the network with 10 hidden layers achieves the highest accuracy and lowest loss when our proposed method is employed. Both initialization methods exhibit lower performance in networks with $3$ layers compared to those with more layers. Moreover, for more complex datasets such as CIFAR-10 and CIFAR-100, the proposed method demonstrated improved performance when training deeper networks.
(a) MNIST
(b) Fashion MNIST
(c) CIFAR10
(d) CIFAR100
Figure 4: Validation accuracy for a tanh FFNN with 50 hidden layers (32 nodes each). Xavier + BN and Xavier + LN represent Xavier initialization with Batch Normalization or Layer Normalization applied every 5 layers, respectively.
Normalization Methods. Xavier initialization is known to cause vanishing gradients and activation problems in deeper networks. These issues can be mitigated by applying Batch Normalization (BN) or Layer Normalization (LN) in the network. Therefore, we compare the proposed method with Xavier, Xavier with BN, and Xavier with LN. To evaluate the effectiveness of normalization, we conducted experiments using a sufficiently deep neural network with $50$ hidden layers. As shown in Figure 4, for datasets with relatively fewer features, such as MNIST and FMNIST, Xavier with normalization converges faster than Xavier. However, for feature-rich datasets such as CIFAR-10 and CIFAR-100, the accuracy of Xavier with normalization is lower than that of Xavier. Normalization typically incurs a $30\%$ computational overhead, and additional hyperparameter tuning is required to determine which layers should apply normalization. In contrast, the proposed method achieves the best performance across all datasets without requiring normalization.
Data Efficiency in Classification Task. Based on the results in Table 2, we evaluated data efficiency on a network with $50$ hidden layers, each containing $64$ nodes, where Xavier exhibited strong performance. As shown in Table 3, the highest average accuracy and lowest average loss over $5$ runs after $100$ epochs are presented for datasets containing $10$, $20$, $30$, $50$, and $100$ samples. The proposed method achieved the best performance across all sample sizes.
@@
Non-uniform Hidden Layer Dimensions. We evaluate the performance of the proposed initialization in networks where hidden layer dimensions are not uniform. As shown in Figure 5, the network consists of 60 hidden layers, where the number of nodes alternates between 32 and 16 in each layer. The proposed method demonstrates improved performance in terms of both loss and accuracy across all epochs on the MNIST and CIFAR-10 datasets. Additionally, Appendix B.2 presents experiments on networks with larger variations in the number of nodes. Motivated by these results, Appendix B.2 further explores autoencoders with significant differences in hidden layer dimensions.
@@
Table 3: Validation accuracy and loss for a 10-layer FFNN (64 nodes per layer) trained on datasets containing 10, 20, 30, 50, and 100 samples. Results show the highest average accuracy and lowest average loss over 5 runs after 100 epochs.
Dataset Method 10 20 30 50 100
Accuracy Loss Accuracy Loss Accuracy Loss Accuracy Loss Accuracy Loss
MNIST Xavier 31.13 2.281 35.03 2.078 45.05 1.771 58.45 1.227 64.02 1.139
Xavier + BN 22.46 2.267 33.73 2.053 37.13 2.042 39.78 1.944 57.51 1.464
Xavier + LN 28.52 2.411 41.54 1.796 41.94 1.886 54.97 1.362 65.11 1.093
Proposed 37.32 2.204 46.79 1.656 48.60 1.645 61.54 1.131 68.44 1.043
FMNIST Xavier 36.16 2.320 41.69 1.814 53.86 1.459 64.53 1.140 63.58 1.048
Xavier + BN 35.44 2.136 38.58 1.925 40.16 1.819 53.93 1.728 59.78 1.237
Xavier + LN 34.94 2.362 37.90 1.793 53.27 1.470 59.50 1.198 62.01 1.073
Proposed 37.31 2.217 49.25 1.651 55.19 1.372 66.14 1.057 67.58 0.914
4.2 Physics-Informed Neural Networks
Xavier initialization is the most commonly employed method for training PINNs (Jin et al., 2021; Son et al., 2023; Yao et al., 2023; Gnanasambandam et al., 2023). In this section, we experimentally demonstrate that the proposed method is more robust across different network sizes and achieves higher data efficiency compared to Xavier initialization with or without normalization methods.
@@ 4.2 Physics-Informed Neural Networks
Experimental Setting. All experiments on Physics-Informed Neural Networks (PINNs) use full-batch training with a learning rate of 0.001. In this section, we solve the Allen-Cahn, Burgers, Diffusion, and Poisson equations using a tanh FFNN-based PINN with 20,000 collocation points. Details on the PDEs are provided in Appendix C.1.
Network Size Independence in PINNs. We construct eight tanh FFNNs, each with $16$ nodes per hidden layer and $5,10,20,30,40,50,60,$ or $80$ hidden layers. As shown in Table 4, for the Allen-Cahn and Burgers’ equations, Xavier+BN and Xavier+LN achieve the lowest loss at a network depth of 30. However, their loss gradually increases as depth grows. In contrast, the proposed method achieves the lowest loss at depths of 50 and 60, respectively, maintaining strong learning performance even in deeper networks. For the Diffusion and Poisson equations, Xavier+LN achieves the lowest loss at depths of 5 and 10, respectively. While all methods exhibit increasing loss as network depth increases, the proposed method consistently maintains lower loss in deeper networks. Similar trends are observed with 32 nodes. Across all tested network sizes and PDEs, the proposed method consistently achieves the lowest loss. The proposed method eliminates the computational overhead and hyperparameter tuning required for normalization methods.
@@
Table 4: A PINN loss is presented for FFNNs with varying numbers of layers $(5,10,20,30,40,50,60,80)$ using the tanh activation function. The top table shows results with $16$ nodes per layer, and the bottom table shows results with $32$ nodes per layer. All models were trained for $300$ iterations using Adam and $300$ iterations using L-BFGS. The median PINN loss at the final iteration for the Burgers, Allen–Cahn, Diffusion, and Poisson equations, computed over 5 runs, is presented.
@@
Diffusion (16 Nodes) 5 10 20 30 40 50 60 80
Xavier 2.52e-03 4.82e-03 9.69e-03 1.33e-02 2.08e-02 1.50e-02 2.92e-02 7.24e-02
Xavier + BN 2.89e-03 5.77e-03 1.05e-02 9.65e-03 2.76e-02 1.07e-02 9.07e-03 1.43e-02
Xavier + LN 1.72e-03 6.10e-03 8.04e-03 9.48e-03 2.14e-02 7.59e-03 2.05e-02 2.21e-02
Proposed 9.14e-04 2.59e-03 2.40e-03 1.01e-03 1.97e-03 1.21e-03 1.12e-03 1.91e-03
Poisson (16 Nodes) 5 10 20 30 40 50 60 80
Xavier 1.52e-02 2.87e-02 1.28e-01 9.82e-02 1.15e-01 1.37e-01 1.82e-01 2.55e-01
Xavier + BN 1.62e-02 2.02e-02 8.72e-02 1.12e-01 2.45e-01 9.85e-02 1.00e-01 1.34e-01
Xavier + LN 5.39e-01 4.40e-02 1.34e-01 3.91 2.52e+02 2.58 9.79e+02 N/A
Proposed 1.37e-02 1.70e-02 4.62e-02 2.43e-02 3.75e-02 4.03e-02 6.07e-02 6.01e-02
@@
Diffusion (32 Nodes) 5 10 20 30 40 50 60 80
Xavier 1.69e-03 6.85e-03 7.63e-03 4.50e-03 8.98e-03 5.67e-03 6.33e-01 1.59
Xavier + BN 1.68e-03 2.66e-03 1.08e-02 6.00e-03 8.58e-03 6.60e-03 5.66e-02 1.69e+02
Xavier + LN 8.16e-04 2.85e-03 8.46e-03 4.57e-03 9.40e-03 1.04e-02 2.42e-01 1.67e+02
Proposed 2.89e-04 8.03e-04 5.25e-04 5.07e-04 5.33e-04 6.17e-04 9.80e-04 1.53e-03
Poisson (32 Nodes) 5 10 20 30 40 50 60 80
Xavier 1.09e-02 1.33e-02 3.13e-02 7.69e-02 6.72e-02 8.90e-02 9.68e+02 1.46e+02
Xavier + BN 1.14e-02 1.47e-02 2.68e-02 3.55e-02 8.25e-02 8.97e-02 4.50e-02 7.75e-01
Xavier + LN 2.36e-02 2.18e-02 3.07e-02 3.85e-01 1.40 4.69 2.60 6.14
Proposed 9.63e-03 8.29e-03 1.41e-02 1.88e-02 1.65e-02 1.85e-02 1.73e-02 3.59e-02
Data Efficiency in PINNs. Based on the results in Table 4, we evaluate data efficiency on a network with $30$ hidden layers, each containing $32$ nodes, where Xavier initialization achieved the lowest PINN loss. As shown in Figure 6, for the Burgers equation, the Mean Absolute Error (MAE) of the proposed initialization differs significantly from that of Xavier initialization across varying numbers of collocation points. In contrast, for the Diffusion equation, the difference in MAE between the two methods becomes more pronounced when the number of collocation points exceeds $20,000$. Additionally, Figure 7 illustrates that increasing the number of collocation points enables PINNs with the proposed initialization to predict solutions with lower absolute error. For detailed experiments on the Burgers equation, please refer to Appendix C.4.
(a) Burgers Equation
(b) Diffusion Equation
Figure 6: Mean absolute error between the exact solution and PINN-predicted solution with varying numbers of collocation points. The FFNN has $30$ hidden layers ($32$ nodes each) and is trained for $300$ iterations using Adam followed by $300$ iterations using L-BFGS. The results are averaged over $5$ experiments.
Figure 7: Absolute error between the exact solution and the PINN-predicted solution for the Diffusion equation with varying numbers of collocation points $(3000,10000,20000,50000)$ using (upper row) Xavier and (lower row) the proposed initialization. The FFNN has 30 hidden layers ($32$ nodes each) and is trained for $300$ iterations using Adam followed by $300$ iterations using L-BFGS. The color bar ranges from 00 to $0.05$, with values outside this range shown in white.